ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 229, 2019

Arendt, W. and I. Chalendar Generators of semigroups on Banach spaces inducing holomorphic semiflows

165

Blomer, V., G. Harcos and P. Maga On the global sup-norm of GL(3) cusp forms

357

Chalendar, I. See W. Arendt Conejero, J. A., C. Lizama, M. Murillo-Arcila and

J. B. Seoane-Sepúlveda Well-posedness for degenerate third order equations with delay and applications to inverse problems

219

Dvornicich, R. and U. Zannier Composite factors of binomials and linear systems in roots of unity

381

Fried, S. and D. Haran Quasi-formations  193

Gallardo-Gutiérrez, E. A. and D. V. Yakubovich On generators of C_{0}-semigroups of composition operators

487

Girão, A., T. Kittipassorn and B. Narayanan Long cycles in Hamiltonian graphs

269

Glücksam, A. Measurably entire functions and their growth  307
Gregory, L. and G. Puninski Ziegler spectra of serial rings  415

Guralnick, R. M. and M. W. Liebeck Permutation representations of nonsplit extensions involving alternating groups

181

Guzmán, O. There is a +-Ramsey MAD family  393

Haran, D. See S. Fried Harcos, G. See V. Blomer Ioppolo, A. and F. Martino Classifying G-graded algebras of ex-

ponent two

341

Jell, P. Tropical Hodge numbers of non-archimedean curves  287
Keller, G. and C. Richard Periods and factors of weak model sets  85

Kittipassorn, T. See A. Girão Liebeck, M. W. See R. M. Guralnick Linial, N., I. Newman, Y. Peled and Y. Rabinovich Extremal

hypercuts and shadows of simplicial complexes

133

Lizama, C. See J. A. Conejero Maga, P. See V. Blomer Martino, F. See A. Ioppolo Merry, W. J. and I. Uljarevic Maximum principles in symplectic

homology

39

Murillo-Arcila, M. See J. A. Conejero Narayanan, B. See A. Girão Newman, I. See N. Linial Ng, T.-W. and C.-F. Wu Nonlinear Loewy factorizable algebraic

ODEs and Hayman’s conjecture

1

Peled, Y. See N. Linial Puninski, G. See L. Gregory Rabinovich, Y. See N. Linial Richard, C. See G. Keller Seoane-Sepúlveda, J. B. See J. A. Conejero Uljarevic, I. See W. J. Merry Viana, M. and J. Yang Continuity of Lyapunov exponents in the

C^{0} topology

461

Wu, C.-F. See T.-W. Ng Yakubovich, D. V. See E. A. Gallardo-Gutiérrez Yang, J. See M. Viana Zannier, U. See R. Dvornicich

Zhao, D. Characters of Iwahori–Hecke algebras  67
Zorin-Kranich, P. A double return times theorem  255